OpenStudy (anonymous):

]Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6.

OpenStudy (geneticrockhopper247):

Okay, so do you know how to find the slope of the equation given (y=mx+b)?

OpenStudy (anonymous):

yes

OpenStudy (geneticrockhopper247):

Okay, so the slope is 3, right?

OpenStudy (geneticrockhopper247):

Also, have you studied point slope form?

OpenStudy (geneticrockhopper247):

If not, it is y-y1=m(x-x1) Or, y-y coordinate given=slope(x-x coordinate given). Because the lines have the same slope, they will be parallel.

OpenStudy (geneticrockhopper247):

@flyguerlkeke You there???

OpenStudy (anonymous):

yes

OpenStudy (geneticrockhopper247):

So does this make sense???

OpenStudy (geneticrockhopper247):

So, if this does make sense, can you take it from here???

OpenStudy (geneticrockhopper247):

Hello???? @flyguerlkeke

OpenStudy (anonymous):

yes

OpenStudy (geneticrockhopper247):

I will take my leave then. Best of luck...

OpenStudy (anonymous):

ok bye